How do you solve showing the steps -2/3(x-1)=36
Answer by unlockmath(1688)

Hello,
First let's get rid of the fraction by multiplying it by 3 to give us:
(3)-2/3(x-1)=36(3) rewritten as:
-2(x-1)=108 Distribute will look like:
-2x+2=108 Subtract 2 from each side.
-2x=106 Divide both sides by -2
x=-53
Plug this into the original equation to see if it works.
RJ Toftness
www.math-unlock.com